OpenStudy (anonymous):

for f(x)=√2x-1, find the following a. the domain of f(x) b. f(2) c. f(x+h)

OpenStudy (***[isuru]***):

hi, for the first part is it \[\sqrt{2x -1} \ or \ \sqrt{2x} -1\] I think it should be the first one ??

OpenStudy (anonymous):

no i need help with like can you find f(x) also find f(2) and find f(x+h)

OpenStudy (***[isuru]***):

yeah !! I can help u , But I need to be sure about the expression first... the expression you've typed .. is it \[f(x) = \sqrt{2x -1}\] or \[f(x) = \sqrt{2x} -1\] the way u have use square root sign, it's not clear !!!

OpenStudy (anonymous):

oh the first one where it covers the whole problem

OpenStudy (***[isuru]***):

got it... and ya want to find f(2) first.... to find f(2) , substitute 2 instead of "x" in f(x) ... i.e -\[f(x) = \sqrt{2x -1}\]\[f(2) = \sqrt{(2\times2) -1}\] now what is the value of f(2) ?

OpenStudy (anonymous):

i'm really bad at math im not sure =(

OpenStudy (***[isuru]***):

nuh.... just forget it... what is the value of this one.. (2x2) -1 ?

OpenStudy (anonymous):

f(2)?

OpenStudy (***[isuru]***):

\[f(2) = \sqrt{(2 \times 2 ) -1}\]\[f(2) = \sqrt{4 -1}\]\[f(2) = \sqrt{3}\] Did ya get it now...? Shall we move to the next part ?

OpenStudy (anonymous):

yes

OpenStudy (***[isuru]***):

so.... the next one asks for f(x + ) to find f(x +h ) u have to substitute "(x + h)" instead of "x" in f(x) i.e \[f(x) = \sqrt{2x -1}\]\[f(x + h ) = \sqrt{2(x + h) - 1}\]\[f(x + h ) = \sqrt{2x +2h- 1}\] that's it !!! did ya get it ?

OpenStudy (anonymous):

yes it's making sense what's the domain of f(x)

OpenStudy (***[isuru]***):

do u know the domain of f(x) ?

OpenStudy (anonymous):

no i was hopin you could help me with that

OpenStudy (anonymous):

do u know it?

OpenStudy (anonymous):

would it be x≥1/2

OpenStudy (***[isuru]***):

domain of f(x) means the all possible values that u can use instead of "x" in f(x).. i.e 2 is a number in the domain of f(x)... because u substitute 2 and u will get a rational answer... but what about -1 ? lets see what will happen to f(-1) \[f(-1) = \sqrt{[2 \times (-1 ) ] -1} \] \[f(-1) = \sqrt{-2-1}\]\[f(-1) =\sqrt{-3}\] Ops!! we've got a negative square root as the answer.. as negative square roots can't be defined rationally .... -1 doesn't belong to the domain of f(x) So.. u have to find the range of value that will give u a positive square root for f(x)... So.. the minimum value for f(x) is 0 ... any value below that will be a negative square root... U will get 0 for f(x) when u find f(1/2)... i.e - when the value of "x" is 1/2 the value of f(x) is 0 the minimum value that f(x) can take is 0 the value of "x" that makes f(x) = 0 is 1/2 so .... the domain of f(x) is \[\left\{ x/x \ \epsilon \ \mathbb{R} ; x \ge 1/2 \right\}\]